OpenStudy (anonymous):

Choose h and k such that the system has (a) no solution (b) a unique solution or (c) many solutions. Give separate answers for each part. x1 - 3x2 = 1 and 2x1 + hx2 = k

OpenStudy (anonymous):

Still trying to get through these review questions from what I missed while ill.

OpenStudy (anonymous):

\[x_1-3x_2=1\\ 2x_1+hx_2=k\]?

OpenStudy (anonymous):

Yep, those are the correct systems. I should have made sure to mention that they were row labels and not exponents. I'm sorry if that was confusing.

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

if you pick some random \(h\) and \(k\) there will probably be one unique solution

OpenStudy (anonymous):

if you multiply the first equation by 2 you get \[2x_1-6x_2=2\\ 2x_2+hx_2=k\]

OpenStudy (anonymous):

if you make \(h=-6\) and \(k=2\) they will be the same line exactly so there will be infinitely many solutions

OpenStudy (anonymous):

if you make \(h=-6\) and \(k=\) anything but 2, they will be parallel lines with no intersection

OpenStudy (anonymous):

Oh wow! How did you figure that out?

OpenStudy (anonymous):

i am the satellite

OpenStudy (anonymous):

ok in truth if you wanted to solve it you would have to match up the coefficients to use "elimination" i just matched up the coefficients of the \(x_1\) term by doubling the first one

OpenStudy (anonymous):

Hehehe, that makes sense.

OpenStudy (anonymous):

Thank you!